OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9)

OpenStudy (anonymous):

@Dianegrace

OpenStudy (anonymous):

The easiest thing to do here is plug in each of the ordered pairs into the equation and see if they work. For example, a) (-6, -14), so plug in -6 for x and -14 for y in your equation 7x-5y=28. So 7(-6)-5(-14)=28 then solve. 7(-6)= -42 and 5(-14)= -70. -42-(-70)= 28. 28=28 and therefore A is an ordered pair that fits the equation. Do the same thing for each of the ordered pairs and you'll have your answer.

OpenStudy (anonymous):

ok thank you ans you that A was one of them?